OpenStudy (anonymous):

What are the solution intervals for |2x – 1| + 6 > 9?

OpenStudy (anonymous):

I 2x-1I = -(2x-1) if x<1/2 = (2x-1) if x>1/2

OpenStudy (anonymous):

so for x<1/2 -(2x – 1) + 6 > 9 and for x>1/2 (2x – 1) + 6 > 9

OpenStudy (anonymous):

solve both inequality separately

OpenStudy (anonymous):

does it make sense?

OpenStudy (anonymous):

You can also do 6-1+9=14 Because we were combining like terms. So we are left with 2x > 14, then we used inverse operations. So we subtract 2 from both sides. So we are left with x>12. So that's the simplified equation.